### v1

# MS-ISSM: Objective Quality Assessment of Point Clouds Using Multi-scale Implicit Structural Similarity

Abstract
The unstructured and irregular nature of point clouds poses a significant challenge for objective quality assessment (PCQA), particularly in establishing accurate perceptual feature correspondence. To tackle this, we propose the Multi-scale Implicit Structural Similarity Measurement (MS-ISSM). Unlike traditional point-to-point matching, MS-ISSM utilizes Radial Basis Functions (RBF) to represent local features continuously, transforming distortion measurement into a comparison of implicit function coefficients. This approach effectively circumvents matching errors inherent in irregular data. Additionally, we propose a ResGrouped-MLP quality assessment network, which robustly maps multi-scale feature differences to perceptual scores. The network architecture departs from traditional flat MLPs by adopting a grouped encoding strategy integrated with Residual Blocks and Channel-wise Attention mechanisms. This hierarchical design allows the model to preserve the distinct physical semantics of luma, chroma, and geometry while adaptively focusing on the most salient distortion features across High, Medium, and Low scales. Experimental results on multiple benchmarks demonstrate that MS-ISSM outperforms state-of-the-art metrics in both reliability and generalization. The source code is available at: https://github.com/ZhangChen2022/MS-ISSM.
I Introduction
Point clouds are fundamental to 3D representation in applications ranging from autonomous driving to augmented and virtual reality (AR/VR) [1, 2]. However, their irregular and unstructured nature makes accurate quality assessment (PCQA) challenging, especially given distortions from noise, sampling, and compression [3]. While subjective assessment provides reliable ground truth, it is costly and time-consuming, necessitating efficient objective metrics that correlate well with human perception [4, 5, 6].
Existing objective PCQA methods generally fall into two categories: projection-based and point-based[7]. Projection-based methods project 3D data onto 2D planes, leveraging mature image quality assessment (IQA) algorithms [8, 9, 10, 11]. However, this dimensionality reduction often causes geometric information loss and introduces viewpoint dependency [12]. Conversely, point-based methods directly utilize 3D spatial features [13, 14]. A common approach involves identifying point-to-point correspondences via nearest-neighbor search to compute geometric or attribute distortions. Yet, due to the unstructured nature of point clouds, establishing accurate correspondence is difficult, and discrete point errors often fail to reflect the continuous surface variations perceived by the human visual system (HVS) [15, 16].
Figure 1: The difference between the MS-ISSM and the traditional point-to-point method.
To address these limitations, we propose the Multi-scale Implicit Structural Similarity Method (MS-ISSM). Building on our previous work utilizing Radial Basis Function (RBF) interpolation [17], we represent point cloud features as implicit functions. Instead of relying on error-prone point-to-point matching, MS-ISSM assesses quality by comparing the coefficients of these implicit functions across matched spatial components. This approach effectively captures surface structural variations and aligns better with human visual perception, as shown in Fig. 1. To handle complex non-linear mappings and incorporate HVS characteristics, we integrate multi-scale analysis and propose a specialized regression network.
The main contributions of this paper are summarized as follows:
  • We propose representing point cloud features using RBF implicit functions. By converting feature differences into implicit function coefficient differences, we mitigate the accumulation of matching errors caused by the irregular nature of point clouds.
  • The proposed coefficient-based comparison eliminates the need for establishing explicit 3D coordinate correspondence between distorted and original point clouds, thereby reducing computational complexity.
  • We design a ResGrouped-MLP quality assessment network. It incorporates a Log-Modulus transformation to handle heavy-tailed feature distributions and integrates Residual Blocks with Channel-wise Attention to adaptively weight physical semantics (luma, chroma, geometry) across multiple scales.
  • Extensive experiments on public datasets demonstrate that MS-ISSM achieves competitive performance compared to state-of-the-art PCQA metrics.
The remainder of this paper is organized as follows: Section II reviews related work. Section III details the problem formulation and theoretical foundation. Section IV describes the proposed MS-ISSM. Section V presents experimental results, and Section VI concludes the paper.
II Related Work
This section reviews existing PCQA methods, categorized into single-scale and multi-scale approaches.
a) Single-scale PCQA Methods: Single-scale methods quantify distortion by measuring geometric or attribute variations between corresponding points. Standard metrics rely on point-to-point Euclidean distances or feature differences, widely adopted in compression standards [7]. Enhancements to these metrics include measuring projection distance along normal directions [13], utilizing Mahalanobis distance to capture spatial distribution [18], or calculating point-to-grid distances [19]. To improve perceptual correlation, other approaches focus on feature disparities, such as angular differences between normal vectors [20] or curvature variations [21]. While these algorithms possess low computational complexity [22], they often fail to align with human visual perception due to the lack of perceptual modeling.
b) Multi-scale PCQA Methods: To better approximate human perception, researchers have integrated multi-scale and joint features. Meynet et al. proposed the Point Cloud Quality Metric (PCQM) [23, 14], a linear combination of curvature, chroma, and brightness. Other hand-crafted feature methods combine geometric statistics with local plane features [24], utilize gradients from local graphs [25, 26], analyze geometric topology alongside color distribution [27], or measure multi-scale spatial potential energy [28], transformational complexity [29] and perception-guided hybrid metrics (PHM) [30]. Additionally, Lazzarotto et al. developed MS-PointSSIM by weighting structural similarity across spatial scales [31].
Recent advancements leverage learning-based frameworks. These include CNN-based mapping of feature differences [32], and GNNs for learning local intrinsic dependencies [33]. Other works employ PCA on local neighborhoods [34] or integrate Spherical Graph Wavelet (SGW) coefficients with Support Vector Regression (FRSVR) [35, 36]. Similarly, Cui et al. combined projected structural similarity with wavelet sub-band features in a learning framework [37]. Wang et al. also explored joint assessment using multi-scale texture features from 2D images and 3D geometric points [38].
Alternatively, projection-based methods evaluate quality by rendering point clouds into 2D images and applying image quality assessment (IQA) models [39, 40, 41]. However, projection alters 3D characteristics, leading to the loss of geometric details such as depth and occlusion relationships. Furthermore, while point-based learning methods show promise, they struggle with the fundamental challenge of establishing accurate point correspondences for distorted data.
III Problem Formulation
III-A Points Correspondence and Perceptual Distortion Measurement
A point cloud is defined as a set of geometric coordinates and associated attributes. Let the original point cloud ${\bf{P}}^{\rm{O}}$} and the distorted point cloud ${\bf{P}}^{\rm{D}}$} be represented as:
[…]
where $\alpha\in\{{\rm{O}},{\rm{D}}\}$ denotes the point cloud type, and $N_{\alpha}$ is the number of points. Each element consists of geometric coordinates ${\bf{p}}_{n}^{\alpha}$ and attributes ${\bf{q}}_{n}^{\alpha}$. Ideally, the perceptual distortion $D({{\bf{P}}^{\rm{O}}},{{\bf{P}}^{\rm{D}}})$ is measured by finding a feature bijection $\psi$ that minimizes the feature difference:
$\begin{array}[]{l}D({{\bf{P}}^{\rm{O}}},{{\bf{P}}^{\rm{D}}})\\ =\mathop{\min}\limits_{\psi:{{\bf{P}}^{\rm{O}}}\to{{\bf{P}}^{\rm{D}}}}\left\{{\frac{1}{{{N_{\psi}}}}\sum\limits_{{\bf{p}}_{i}^{\rm{O}}\in{{\bf{P}}^{\rm{O}}}}{{{\left\|{{{\rm{M}}_{\rm{O}}}({\bf{p}}_{i}^{\rm{O}})-{{\rm{M}}_{\rm{D}}}\left({\psi\left({{\bf{p}}_{i}^{\rm{O}}}\right)}\right)}\right\|}_{2}}}}\right\}\end{array}\text{,}$ (2)
where ${{\rm{M}}_{\alpha}}(\cdot)$ extracts features (e.g., geometry, color). However, since $N_{\rm{O}}$} often differs from $N_{\rm{D}}$}, a strict bijection is impractical. Consequently, classical methods approximate this using nearest-neighbor search to compute the symmetric distortion:
[…]
where $\varphi_{{\rm{O}}\to{\rm{D}}}$} denotes the injective mapping determined by nearest-neighbor search. Due to the unordered nature of point cloud data and the random distribution of points in space, a simple nearest-neighbor search may result in incorrect mapping. For example, if the points in ${\bf{P}}^{O}$} are denser than those in ${\bf{P}}^{D}$} or if there is a spatial distribution bias between the two, the nearest-neighbor search leads to inaccurate matches. This, in turn, would affect the calculation of feature differences and, ultimately, the distortion measurement, as shown in Fig.1. To achieve accurate feature correspondence, in our earlier work[17], we obtained a bijective set of point features by using a feature interpolation function. However, this method only utilized single-scale luminance values. Additionally, the distortion calculation method based on points struggles to account for changes in the local structure of the point cloud, leading to distortion results that differ from actual perception.
Figure 2: The schematic diagram depicts implementing the MS-ISSM solution. (1) Multi-scale features are extracted from the normalized distorted and original point clouds. The chroma, luma, and curvature features of each point cloud are calculated under high-, medium-, and low-quality conditions. (2) The RBF implicit representation is used to calculate the coefficient values for each feature, and multi-scale feature coefficient differences are calculated. (3) The ResGrouped-MLP is designed to map the multi-scale features coefficient differences to perceptual quality scores.
III-B Feature Implicit Representation and Multiscale Perceptual Distortion Calculation
To address the above challenges, considering that exact point-to-point correspondence is difficult to achieve, for a local space, we map the feature set of the point cloud to a feature function, as shown in:
$\left\{\begin{array}[]{l}\left\{{{\bf{p}}_{i}^{\rm{O}},{{\rm{M}}_{\rm{O}}}\left({{\bf{p}}_{i}^{\rm{O}}}\right)}\right\}_{i=1}^{{N_{\rm{O}}}}\to{f^{\rm{O}}}\left({{\bf{p}}_{i}^{\rm{O}}{\rm{,}}{{\bf{W}}^{\rm{O}}}}\right)={{\rm{M}}_{\rm{O}}}\left({{\bf{p}}_{i}^{\rm{O}}}\right)\\ \left\{{{\bf{p}}_{j}^{\rm{D}},{{\rm{M}}_{\rm{D}}}\left({{\bf{p}}_{j}^{\rm{D}}}\right)}\right\}_{j=1}^{{N_{\rm{D}}}}\to{f^{\rm{D}}}\left({{\bf{p}}_{j}^{\rm{D}}{\rm{,}}{{\bf{W}}^{\rm{D}}}}\right)={{\rm{M}}_{\rm{D}}}\left({{\bf{p}}_{j}^{\rm{D}}}\right)\end{array}\right$$\text{,}$ (4)
where ${f^{\rm{O}}}\left({{\bf{p}}_{i}^{\rm{O}}{\rm{,}}{{\bf{W}}^{\rm{O}}}}\right)$ and ${f^{\rm{D}}}\left({{\bf{p}}_{j}^{\rm{D}}{\rm{,}}{{\bf{W}}^{\rm{D}}}}\right)$ represent the feature functions of the original and distorted point clouds, respectively. The matrices ${{\bf{W}}^{\rm{O}}}=\left\{{{w_{k}}^{\rm{O}}}\right\}_{k=1}^{K}$ and ${{\bf{W}}^{\rm{D}}}=\left\{{{w_{k}}^{\rm{D}}}\right\}_{k=1}^{K}$ are the coefficient matrices of the implicit functions of ${\bf{P}}^{\rm{O}}$} and ${\bf{P}}^{\rm{D}}$}, respectively. This approach transforms the feature difference into the error between their corresponding feature functions, as shown in Fig. 1.
Since ${{\bf{W}}^{\rm{O}}}=\left\{{{w_{k}}^{\rm{O}}}\right\}_{k=1}^{K}$ and ${{\bf{W}}^{\rm{D}}}=\left\{{{w_{k}}^{\rm{D}}}\right\}_{k=1}^{K}$ have the same functional form, we calculate the feature function difference by the implicit function coefficients, as shown in:
$\left\{\begin{array}[]{l}D^{\prime}({{\bf{P}}^{\rm{O}}},{{\bf{P}}^{\rm{D}}})=g\left({{{d^{\prime}}_{1}},{{d^{\prime}}_{2}},...,{{d^{\prime}}_{k}},...,{{d^{\prime}}_{K}}}\right)\\ {{d^{\prime}}_{k}}=\left|{\frac{{w_{k}^{\rm{O}}-w_{k}^{\rm{D}}}}{{\max\left\{{w_{k}^{\rm{O}}{\rm{,}}w_{k}^{\rm{D}}}\right\}}}}\right|\end{array}\right$$\text{,}$ (5)
where $d_{k}$} represents the difference between the individual function coefficients, and $g()$ is the nonlinear mapping from the feature function coefficient difference to the perceptual difference, and this mapping is obtained through a regression model. This method addresses the difficulty of point-to-point correspondence matching while considering the structural changes in local features.
Furthermore, considering the multi-scale nature of human visual perception, we incorporate the differences at low, medium, and high scales into the final distortion calculation. The final distortion is expressed as:
${D_{pro}}=g\left({\underbrace{{{d^{\prime}}_{1}},...,{{d^{\prime}}_{K}}}_{\rm{L}},\underbrace{{{d^{\prime}}_{1}},...,{{d^{\prime}}_{K}}}_{\rm{M}},\underbrace{{{d^{\prime}}_{1}},...,{{d^{\prime}}_{K}}}_{\rm{H}}}\right)\text{.}$ (6)
Figure 3: Schematic illustration of the Multi-scale implicit feature extraction framework for point cloud.
IV Proposed MS-ISSM
In this section, we first present the general framework of the proposed MS-ISSM in subsection $A$. Then, the implementation details of each module are described in subsections BDB-D, respectively.
IV-A Overview
Due to the complexity of the human visual system (HVS), extracting features from point clouds and mapping them to precise perceptual quality metrics is a challenging task[22]. To address this issue and streamline the computation process, we approach it in three steps.
In the first step, we extract multi-scale features from the normalized distorted and original point clouds. We focus on the chroma, luma, and curvature features of each point cloud under high-, medium-, and low-quality conditions.
In the second step, we apply the RBF implicit representation to the spatial scale features of the obtained distorted point cloud and the original point cloud, calculating the coefficient values for each feature’s implicit representation.
In the third step, we propose the ResGrouped-MLP to map the multi-scale feature coefficient differences to perceptual quality scores.
The overall process of the proposed MS-ISSM is illustrated in Fig. 2.
IV-B Multi-scale Features Extraction
To capture complex perceptual changes, we utilize three physical features that align with the HVS: curvature, luma, and chroma. Curvature describes the local surface geometry, reflecting sensitivity to both fine details and global structure [21]. Luma and chroma, calculated from the point cloud’s color components [7], represent light intensity and color distribution, respectively.
To ensure generalization across varying geometric scales, we normalize the geometric components of both distorted and original point clouds as follows:
${\bf{\hat{p}}}_{n}^{\alpha}=\frac{1024\cdot({\bf{p}}_{n}^{\alpha}-{\bf{p}}_{min})}{L_{max}}\text{,}$ (7)
where ${\bf{\hat{p}}}_{n}^{\alpha}$ denotes the normalized coordinates. $L_{max}$ is the maximum edge length of the bounding box of the original point cloud ${\bf{P}}^{O}$}, and ${\bf{p}}_{\min}$} is the coordinate-wise minimum vector. The resulting normalized point cloud is denoted as ${{\bf{\hat{P}}}^{\alpha}}=\left\{{{\bf{\hat{p}}}_{n}^{\alpha},{\bf{q}}_{n}^{\alpha}}\right\}$.
Furthermore, inspired by the multi-layered perceptual mechanism of human vision, multi-scale point clouds (high, medium, and low) are generated from both the original and distorted inputs through voxel grid downsampling. The voxel sizes are set to 2.0, 4.0, and 8.0, respectively. These values follow a dyadic progression, consistent with the hierarchical octree decomposition widely used in point cloud compression standards [7]. This geometric progression allows for a systematic separation of spatial frequency components: the smallest scale captures high-frequency details (e.g., texture and noise), while the largest scale retains low-frequency structural information (e.g., global shape), ensuring a comprehensive evaluation of perceptual quality. Mimicking the way human vision adapts to different environments, this method focuses on various levels of detail, as shown in Fig. 3.
Figure 4: The Proposed ResGrouped-MLP Network.
IV-C Feature Implicit Representation
Based on the attention mechanism of HVS, we measure perceptual distortion by comparing the feature differences of local point clouds, rather than individual points. The features of the local point cloud are implicitly represented using RBF [42]:
$\begin{array}[]{l}f_{\bf{F}}^{\alpha,\beta}({\bf{\hat{p}}}^{\alpha,\beta})\\ =\eta_{\bf{F}}^{\alpha,\beta}({\bf{\hat{p}}}^{\alpha,\beta})+\sum\limits_{n=1}^{N_{\beta}^{\alpha}}{\left[{\omega_{{\bf{F}},n}^{\alpha,\beta}\cdot{\phi^{\alpha,\beta}}\left({{{\left\|{{\bf{\hat{p}}}^{\alpha,\beta}-{\bf{\hat{p}}}_{n}^{\alpha,\beta}}\right\|}_{2}}}\right)}\right]}\end{array}\text{,}$ (8)
where $\beta\in\{{\rm{H}},{\rm{M}},{\rm{L}}\}$ represents different spatial scales: high, medium, and low. ${\bf{F}}\in\{\rm{Cu,Y,Cr}\}$ denotes feature types: curvature, luma, and chroma, and $N^{\alpha,\beta}$} indicates the number of points that influence the implicit function. $f_{\bf{F}}^{\alpha,\beta}({\bf{\hat{p}}}_{n}^{\alpha,\beta})$ corresponds to the implicit function associated with the current feature ${\bf{F}}^{\alpha,\beta}$}, and ${\bf{\hat{p}}}^{\alpha,\beta}\in{{\bf{\hat{p}}}^{\alpha,\beta}}$. $\omega_{{\bf{F}},n}^{\alpha,\beta}$ denote the weight coefficients. $\eta_{\bf{F}}^{\alpha,\beta}({\bf{\hat{p}}}^{\alpha,\beta})$ is a three-variable polynomial with a maximum degree of 3. It is commonly expressed as
$\eta_{\bf{F}}^{\alpha,\beta}({\bf{\hat{p}}}^{\alpha,\beta})={\rm{a}}_{\bf{F}}^{\alpha,\beta}\cdot\hat{x}^{\alpha,\beta}+{\rm{b}}_{\bf{F}}^{\alpha,\beta}\cdot\hat{y}^{\alpha,\beta}+{\rm{c}}_{\bf{F}}^{\alpha,\beta}\cdot\hat{z}^{\alpha,\beta}+{\rm{d}}_{\bf{F}}^{\alpha,\beta}\text{,}$ (9)
where ${\rm{a}}_{\bf{F}}^{\alpha,\beta}{\rm{,b}}_{\bf{F}}^{\alpha,\beta}{\rm{,c}}_{\bf{F}}^{\alpha,\beta}{\rm{,d}}_{\bf{F}}^{\alpha,\beta}$ are constant coefficients. $\hat{x}^{\alpha,\beta}$, $\hat{y}^{\alpha,\beta}$, and $\hat{z}^{\alpha,\beta}$ represent coordinates of point ${{\bf{\hat{p}}}}^{\alpha,\beta}$} at $x,y,z$ directions, respectively. To ensure orthogonality, the weight coefficients $\omega_{{\bf{F}},n}^{\alpha,\beta}$ must satisfy the following constraint conditions:
$\sum\limits_{n=1}^{{N^{\alpha,\beta}}}{\omega_{{\bf{F}},n}^{\alpha,\beta}}=\sum\limits_{n=1}^{{N^{\alpha,\beta}}}{\omega_{{\bf{F}},n}^{\alpha,\beta}\hat{x}_{n}^{\alpha,\beta}}=\sum\limits_{n=1}^{{N^{\alpha,\beta}}}{\omega_{{\bf{F}},n}^{\alpha,\beta}\hat{y}_{n}^{\alpha,\beta}}=\sum\limits_{n=1}^{{N^{\alpha,\beta}}}{\omega_{{\bf{F}},n}^{\alpha,\beta}\hat{z}_{n}^{\alpha,\beta}}=0$
And by inputting all coordinates of ${{\bf{\hat{p}}}}^{\alpha,\beta}$} in ${\bf{\hat{P}}}^{\alpha,\beta}$} into Eq. (8), we determine the coefficients of $\eta_{\bf{F}}^{\alpha,\beta}({\bf{\hat{p}}}^{\alpha,\beta})$ and $\omega_{{\bf{F}},n}^{\alpha,\beta}$ through the following equation:
${\mathbf{X}^{\alpha,\beta}}\cdot\mathbf{W}_{\bf{F}}^{\alpha,\beta}={\mathbf{Y}_{\bf{F}}^{\alpha,\beta}}\text{.}$ (11)
In Eq. (11), $\mathbf{Y}_{\bf{F}}^{\alpha,\beta}$ is the feature matrix, as shown in
[…]
${\mathbf{X}}^{\alpha,\beta}$} is the coordinate matrix, as represented in
[…]
where $\phi_{12}^{\alpha,\beta}$} is equal to $\phi^{\alpha,\beta}\left({{{\left\|{{{\bf{\hat{p}}}_{1}^{\alpha,\beta}}-{{\bf{\hat{p}}}_{2}^{\alpha,\beta}}}\right\|}_{2}}}\right)$}. And $\mathbf{W}_{\bf{F}}^{\alpha,\beta}$ is the weight matrix, as represented in
[…]
Since the number of coefficients in the weight matrix ${\mathbf{W}}_{\bf{F}}^{\alpha,\beta}$ is determined by the number of points $N^{\alpha,\beta}$}, to simplify the computational process and facilitate the comparison of distortions using the coefficients, we downsample $\bf{\hat{P}}^{O}$ to obtain a set of reference points, denoted as ${{\bf{P}}^{\rm{R}}}=\left\{{{\bf{p}}_{t}^{\rm{R}}}\right\}_{t=1}^{{N_{\rm{R}}}}$. Using nearest-neighbor search, for each point ${\bf{p}}_{t}^{\rm{R}}$ in the reference point set, we find the 30 closest neighbors in both $\left\{{{\bf{\hat{p}}}_{i}^{\rm{O}}}\right\}_{i=1}^{{N_{\rm{O}}}}$ and $\left\{{{\bf{\hat{p}}}_{j}^{\rm{D}}}\right\}_{j=1}^{{N_{\rm{D}}}}$, which are then used to compute the implicit function for the region around each reference point. Finally, the features of ${\bf{\hat{P}}}^{O}$ and ${\bf{\hat{P}}}^{D}$ can be represented as tensors $\mathbf{C}^{\alpha}$}:
[…]
By substituting $\mathbf{C}^{\rm{O}}$ and $\mathbf{C}^{\rm{D}}$ into Eq. (5) and (6), the quality score of ${\bf{P}}^{\rm{D}}$ is calculated. Considering the simplicity of the algorithm and to avoid an excessive number of comparison coefficients, we take the average of weight coefficients differences. And the function $g()$ is obtained through the ResGrouped-MLP network as follow.
IV-D ResGrouped-MLP Regression
We propose the ResGrouped-MLP, a hierarchical deep learning framework designed to robustly map hand-crafted point cloud features to subjective quality scores (MOS). Addressing the limitations of flat networks, we adopt a ”Split-Transform-Merge” architecture to preserve the distinct physical semantics of features. The framework integrates a novel Log-Modulus preprocessing and a multi-scale attention mechanism, as shown in Fig. 4.
Figure 5: Data Pipeline and Validation Strategy
Log-Modulus Preprocessing: Statistical features extracted from point clouds often exhibit a heavy-tailed distribution, where standard Z-score normalization can lead to gradient instability and poor convergence. To address this, we introduce the Log-Modulus transformation before normalization. Unlike simple logarithmic transforms, this method handles both positive and negative values while compressing the dynamic range:
$\tilde{\rm{x}}=\text{sign}(\rm{x})\cdot\ln(1+|\rm{x}|)\text{,}$ (16)
where $\rm{x}$ denotes the raw feature value. This transformation effectively rectifies the data distribution towards a quasi-normal form, allowing the subsequent network to focus on underlying feature patterns rather than being biased by outliers, as shown in Fig. 5.
The ResGrouped-MLP Architecture: Point cloud quality perception relies on the interplay between feature scales (High, Medium, Low) and attribute channels (Luma, Chroma, Geometry). Simply concatenating these features risks losing their distinct physical meanings. Therefore, we design a hierarchical architecture composed of three key stages:
a) Deep Grouped Encoders: Instead of a generic fully connected layer, we treat each scale-channel pair as an independent group. We employ Residual Blocks for feature encoding, formulated as $x+\mathcal{F}(x)$, where $\mathcal{F}$ represents a dual-layer perceptron with Batch Normalization and SiLU activation. This grouped design isolates the specific distortion characteristics of each channel, preventing information interference at early stages while mitigating the vanishing gradient problem.
b) Scale-wise Attention Fusion: Since different color channels contribute unequally to human perception, we introduce a Channel Attention Block to recalibrate features within each scale. The mechanism adaptively learns weights to highlight salient distortions:
$\mathbf{F}{scale}^{\prime}=\mathbf{F}{scale}\otimes\sigma(\text{MLP}(\mathbf{F}{scale}))\text{,}$ (17)
where $\mathbf{F}{scale}$ is the concatenated feature vector, $\otimes$ denotes element-wise multiplication, and $\sigma(\cdot)$ is the Sigmoid function. The internal MLP utilizes a bottleneck structure with a reduction ratio $r=4$. This bottleneck compresses the feature space to aggregate global information before restoring it, enabling the network to dynamically suppress noise and emphasize the most relevant feature channels.
c) Global Hierarchical Regression: Finally, the refined features from H, M, and L scales are concatenated and passed through a Global Attention module. This ensures the model captures the complex interaction between global geometry (Low scale) and local fine-grained details (High scale) before mapping them to the final quality score.
Loss Function: To ensure prediction accuracy, linearity, and monotonic consistency, we utilize a hybrid loss function combining Mean Squared Error (MSE), Pearson Linear Correlation Coefficient (PLCC) loss, and Margin Ranking loss:
$\mathcal{L}{total}=\mathcal{L}{MSE}+\lambda_{1}\mathcal{L}{PLCC}+\lambda_{2}\mathcal{L}{Rank}\text{,}$ (18)
where $\lambda_{1}$ and $\lambda_{2}$ are weighting parameters used to balance the optimization objectives.
Implementation and Validation: The model is trained using the AdamW optimizer with a weight decay of $10^{-2}$ and a Cosine Annealing scheduler for 80 epochs (batch size 32). For validation, we employ a rigorous Repeated Random Shuffle Split strategy (5 rounds), partitioning the dataset into 60% training and 40% testing sets. This large test ratio ensures the model is evaluated on a substantial amount of unseen data, demonstrating its generalization capability.
V Experimental Evaluations
This section uses four publicly available point cloud subjective datasets to validate the proposed method’s effectiveness in perceptual evaluation. We compare the quality assessment results of our method with those of classic and state-of-the-art (SOTA) PCQA metrics. By analyzing the results across various datasets, we can assess the robustness and generalization of our method in different distortions.
TABLE I: Performance comparison of different PCQA metrics. The best and second-best are highlighted in red and blue, respectively.
Method Crit. Datasets M-PCCD [43] WPC [40] ICIP [44] SJTU [10] ALL PSNR-p2p [7] PLCC 0.408 0.318 0.566 0.592 0.404 SROCC 0.475 0.297 0.579 0.662 0.398 KROCC 0.354 0.202 0.435 0.493 0.278 RMSE 0.446 0.448 0.331 0.376 0.504 PSNR-p2pl [7] PLCC 0.422 0.250 0.560 0.534 0.391 SROCC 0.489 0.229 0.583 0.602 0.380 KROCC 0.368 0.159 0.458 0.444 0.269 RMSE 0.446 0.457 0.342 0.400 0.504 PSNR-YUV [7] PLCC 0.603 0.274 0.769 0.535 0.439 SROCC 0.620 0.260 0.777 0.553 0.414 KROCC 0.442 0.177 0.580 0.385 0.280 RMSE 0.313 0.353 0.212 0.363 0.396 GraphSIM [25] PLCC 0.912 0.597 0.889 0.854 0.716 SROCC 0.936 0.556 0.888 0.850 0.704 KROCC 0.786 0.412 0.723 0.649 0.524 RMSE 0.246 0.223 0.193 0.194 0.205 MS-GraphSIM [26] PLCC 0.893 0.614 0.892 0.888 0.737 SROCC 0.922 0.581 0.890 0.882 0.726 KROCC 0.761 0.453 0.727 0.692 0.560 RMSE 0.250 0.224 0.197 0.200 0.208 PCQM [14] PLCC 0.913 0.310 0.869 0.576 0.467 SROCC 0.916 0.378 0.956 0.772 0.606 KROCC 0.748 0.306 0.830 0.589 0.470 RMSE 0.288 0.241 0.229 0.243 0.242 PSSIM [27] PLCC 0.592 0.497 0.675 0.759 0.595 SROCC 0.716 0.424 0.778 0.721 0.579 KROCC 0.538 0.300 0.601 0.525 0.414 RMSE 0.262 0.227 0.200 0.207 0.216 MS-PSSIM [31] PLCC 0.814 0.658 0.839 0.876 0.703 SROCC 0.927 0.655 0.841 0.900 0.705 KROCC 0.774 0.478 0.669 0.721 0.508 RMSE 0.264 0.224 0.212 0.208 0.221 TCDM [29] PLCC 0.921 0.795 0.958 0.924 0.847 SROCC 0.947 0.807 0.964 0.924 0.861 KROCC 0.808 0.610 0.863 0.760 0.664 RMSE 0.249 0.211 0.262 0.190 0.204 FRSVR [36] PLCC 0.754 0.243 0.772 0.549 0.438 SROCC 0.818 0.274 0.761 0.679 0.480 KROCC 0.621 0.175 0.587 0.478 0.317 RMSE 0.206 0.276 0.160 0.198 0.221 PHM [30] PLCC 0.935 0.813 0.831 0.892 0.796 SROCC 0.958 0.849 0.953 0.896 0.852 KROCC 0.829 0.660 0.833 0.717 0.656 RMSE 0.275 0.238 0.211 0.216 0.227 MS-ISSM (Ours) PLCC 0.944 0.931 0.925 0.886 0.913 SROCC 0.949 0.935 0.953 0.863 0.914 KROCC 0.811 0.774 0.833 0.701 0.751 RMSE 0.224 0.203 0.193 0.214 0.188
TABLE II: Performance comparison of classic and SOTA metrics on different distortion types. The best and second-best are highlighted in red and blue, respectively.
Distortion types
Metrics Criterias Ds Ns Oc Mx Tc Vc Rank
PLCC 0.275 0.589 0.412 0.633 0.426 0.207 10.17
SROCC 0.339 0.609 0.324 0.646 0.437 0.293 10.33
KROCC 0.233 0.442 0.229 0.468 0.307 0.199 10.50
PSNR-p2p RMSE 0.411 0.308 0.322 0.288 0.345 0.326 11.17
PLCC 0.208 0.607 0.421 0.635 0.440 0.240 9.33
SROCC 0.180 0.622 0.326 0.598 0.435 0.334 10.33
KROCC 0.122 0.453 0.230 0.425 0.311 0.230 10.17
PSNR-p2pl RMSE 0.440 0.306 0.318 0.296 0.324 0.323 10.83
PLCC 0.655 0.476 0.567 0.604 0.276 0.452 10.00
SROCC 0.646 0.507 0.613 0.680 0.271 0.448 10.00
KROCC 0.443 0.368 0.419 0.522 0.192 0.310 10.17
PSNR-YUV RMSE 0.287 0.310 0.317 0.299 0.381 0.230 10.83
PLCC 0.939 0.888 0.856 0.859 0.494 0.630 4.50
SROCC 0.891 0.894 0.862 0.850 0.442 0.603 6.00
KROCC 0.675 0.710 0.678 0.640 0.328 0.425 5.83
GraphSIM RMSE 0.247 0.173 0.214 0.204 0.243 0.167 3.00
PLCC 0.935 0.907 0.882 0.903 0.507 0.652 3.33
SROCC 0.903 0.910 0.882 0.869 0.455 0.640 4.33
KROCC 0.712 0.733 0.699 0.667 0.337 0.452 4.50
MS-GraphSIM RMSE 0.254 0.177 0.217 0.209 0.243 0.169 4.33
PLCC 0.120 0.877 0.607 0.429 0.293 0.594 9.17
SROCC 0.159 0.916 0.896 0.852 0.387 0.568 7.00
KROCC 0.036 0.743 0.707 0.667 0.326 0.407 6.33
PCQM RMSE 0.291 0.218 0.255 0.249 0.262 0.182 9.17
PLCC 0.697 0.629 0.633 0.738 0.218 0.234 9.00
SROCC 0.719 0.694 0.729 0.894 0.290 0.358 8.17
KROCC 0.563 0.514 0.524 0.719 0.215 0.246 8.17
PSSIM RMSE 0.286 0.212 0.246 0.240 0.260 0.179 8.00
PLCC 0.844 0.866 0.798 0.893 0.654 0.551 5.83
SROCC 0.874 0.879 0.839 0.911 0.627 0.544 5.33
KROCC 0.672 0.683 0.647 0.737 0.444 0.378 5.50
MS-PSSIM RMSE 0.278 0.185 0.234 0.220 0.235 0.169 5.67
PLCC 0.898 0.909 0.899 0.932 0.834 0.647 2.67
SROCC 0.873 0.906 0.901 0.935 0.845 0.657 3.33
KROCC 0.671 0.732 0.723 0.779 0.654 0.477 3.17
TCDM RMSE 0.261 0.170 0.219 0.205 0.222 0.165 3.00
PLCC 0.809 0.522 0.733 0.560 0.408 0.417 9.00
SROCC 0.644 0.611 0.752 0.594 0.449 0.453 9.00
KROCC 0.465 0.432 0.552 0.404 0.292 0.319 9.50
FRSVR RMSE 0.265 0.178 0.180 0.211 0.235 0.164 3.50
PLCC 0.906 0.913 0.824 0.896 0.763 0.663 3.17
SROCC 0.897 0.916 0.910 0.909 0.790 0.664 2.33
KROCC 0.693 0.749 0.735 0.742 0.599 0.468 2.50
PHM RMSE 0.277 0.195 0.236 0.229 0.246 0.173 6.83
PLCC 0.900 0.919 0.892 0.920 0.930 0.914 1.83
SROCC 0.934 0.924 0.905 0.886 0.922 0.913 1.83
KROCC 0.737 0.756 0.737 0.704 0.757 0.754 1.67
MS-ISSM (Ours) RMSE 0.228 0.166 0.214 0.208 0.216 0.160 1.67
V-A Datasets and PCQA Metrics under Comparison
To verify the performance of the proposed method across different types of distortions, we use four point cloud subjective datasets for assessment. These datasets include: SJTU [10], WPC [40], M-PCCD[43], and ICIP [44]. We also combine these four datasets into a comprehensive dataset, ALL, to validate the stability and reliability of objective metrics. It is important to note that the subjective rating scales differ across these datasets. To address this, we map the subjective scores from all four datasets onto a common scale of [0, 1] for consistency. In addition, this paper compares the proposed algorithm with 11 classic and SOTA PCQA metrics.
V-B Evaluation Criteria
To ensure alignment between the subjective ratings and objective predictions of different metrics, we standardize the objective predictions to a consistent dynamic range based on guidance from the Video Quality Expert Group (VQEG) [45]. Subsequently, we use Pearson’s linear correlation coefficient (PLCC), Spearman’s rank order correlation coefficient (SROCC), Kendall’s rank order correlation coefficient (KROCC), and root mean square error (RMSE) to evaluate the performance of various metrics, representing their linearity, monotonicity, and accuracy, respectively. Higher PLCC, SROCC, and KROCC values indicate superior metric performance, while lower RMSE values suggest better accuracy. To normalize the scores of objective quality assessment metrics onto a uniform scale, we apply the logistic regression method recommended by VQEG.
V-C Performance Comparison
We evaluate the performance of various PCQA metrics using different datasets. The overall evaluation results for each PCQA metric across these datasets are presented in Table I. To facilitate direct comparison, the best and second-best performing metrics for each dataset are highlighted in red and blue, respectively. Specifically, the proposed method ranks highest on the WPC dataset compared to the other 11 PCQA metrics. On the M-PCCD dataset, the proposed method achieved first place in PLCC, and second place in SROCC, KROCC, and RMSE. On the ICIP dataset, it performs second only to TCDM. Although the proposed method does not achieve the best performance on the SJTU dataset, it closely trails the top-performing methods. On the SJTU dataset, MS-ISSM yields results of [0.886, 0.863, 0.701, 0.214], compared with the best-performing method, TCDM, which achieves [0.924, 0.924, 0.760, 0.190]. Notably, on the combined ALL dataset comprising all four datasets, the proposed method demonstrates superior performance with [0.913, 0.914, 0.751, 0.188], closely followed by TCDM with [0.847, 0.861, 0.664, 0.204]. Overall, the proposed method demonstrates superior performance.
Table II presents the performance of the proposed method and other metrics across various distortion types, including octree-based compression distortion (Oc), video-based compression distortion (Vc), trisoup-based compression distortion (Tc), noise distortion (Ns), downsampling distortion (Ds), and mixing distortion (Mix). As with previous tables, the best and second-best-performing metrics for each distortion type are highlighted in red and blue, respectively. In general, the proposed MS-ISSM exhibits dominating performance in Downsampling, Noise, Trisoup-based compression, and Video-based compression distortions. Specifically, it achieves the top rank across all four indicators (PLCC, SROCC, KROCC, and RMSE) for these categories. For Octree-based compression, MS-ISSM secures the best KROCC score, with other metrics closely following the top performer.
Regarding Mixing distortion, MS-ISSM performs slightly lower than TCDM. This is primarily because the mixing distortion samples originate entirely from the SJTU dataset, on which TCDM’s parameters were explicitly fitted, granting it a distributional advantage in this specific scenario. However, in the more challenging Video-based compression scenario, MS-ISSM achieves a significant lead with performance scores of [0.914, 0.913, 0.754, 0.160], far surpassing the second-best method, which only reaches [0.663, 0.664, 0.477, 0.164].
When ranking all methods across different distortion types, the proposed method achieves the best overall performance in terms of PLCC, SROCC, KROCC, and RMSE. This generalization capability confirms the effectiveness of our design: (1) Robustness of Implicit Surface Reconstruction: The RBF-based implicit representation reconstructs continuous surfaces to resolve sparsity and geometric jitter. This bypasses discrete point-to-point matching errors, yielding superior stability in V-PCC and compression distortions. (2) Comprehensiveness of Multi-scale Strategy: Our hierarchical approach captures both global structural shifts (Low Scale) from downsampling and local high-frequency artifacts (High Scale) from compression. (3) Adaptive Non-linear Mapping: The ResGrouped-MLP combines Log-Modulus transformation for distribution rectification with Channel-wise Attention for adaptive feature weighting, ensuring robust prediction in mixed distortion scenarios.
Additionally, to further compare the performance of the MS-ISSM method and the RBFIM method in terms of compression distortion, we evaluated datasets containing compression distortion from the three aforementioned datasets. Based on the type of compression distortion, the datasets were divided into G-PCC compression distortion and V-PCC compression distortion. The experimental results in Table III reveal that RBFIM performs better on the ICIP G-PCC compression distortion portion. However, MS-ISSM significantly outperforms RBFIM in the V-PCC portion of both M-PCCD and WPC, and performs better across all compression distortion types. This indicates that the multi-scale implicit feature method exhibits stronger generalization and robustness.
TABLE III: Comparison of the RBFIM and MS-ISSM methods under different compression distortions.
Datasets criteria RBFIM[17] MS-ISSM
ICIP-GPCC PLCC 0.993 0.835
SROCC 0.971 0.841
KROCC 0.870 0.610
RMSE 0.020 0.194
ICIP-VPCC PLCC 0.969 0.965
SROCC 0.976 0.996
KROCC 0.895 0.995
RMSE 0.067 0.130
WPC-GPCC PLCC 0.841 0.864
SROCC 0.847 0.866
KROCC 0.655 0.731
RMSE 0.193 0.153
WPC-VPCC PLCC 0.482 0.923
SROCC 0.473 0.921
KROCC 0.343 0.771
RMSE 0.190 0.171
MPCCD-GPCC PLCC 0.734 0.935
SROCC 0.673 0.924
KROCC 0.544 0.842
RMSE 0.280 0.216
MPCCD-VPCC PLCC 0.489 0.838
SROCC 0.460 0.801
KROCC 0.288 0.632
RMSE 0.245 0.190
ALL PLCC 0.611 0.886
SROCC 0.566 0.875
KROCC 0.406 0.712
RMSE 0.259 0.172
We compare the average running times across four datasets on an Intel Core i7-8809G CPU @3.10GHz. As shown in Fig. 6, our method achieves superior efficiency, surpassed only by FRSVR. This speed advantage stems from the proposed implicit feature representation, which eliminates the computationally expensive point-to-point matching process. Furthermore, the pre-computed features of the reference point cloud can be reused across different distortion types, avoiding redundant computation. Conversely, GraphSIM and MS-GraphSIM incur high costs due to keypoint sampling and graph construction, while MS-PSSIM is burdened by high-dimensional multi-scale processing. Although simple single-scale metrics like p2p remain computationally light, our method offers a better trade-off between processing speed and multi-scale performance, making it highly suitable for large-scale PCQA.
Figure 6: Time complexity of different PCQA methods on ALL datasets.
The synthesis of all test results indicates that the proposed MS-ISSM successfully aligns the distorted point cloud with the original using implicit structure by corresponding to the feature function coefficient. This alignment is important as it enhances the correspondence of features within point clouds, facilitating a more precise measurement of distortion. By improving feature correspondence, the MS-ISSM aligns and preserves point clouds’ intrinsic geometric and topological attributes, which are critical for accurate distortion measurement.
Figure 7: Performance comparison among different feature types.
V-D Ablation Studies
To comprehensively validate the effectiveness of the proposed MS-ISSM framework, we conducted extensive ablation studies. These experiments are designed to investigate the contribution of three key aspects: 1. the multi-modal implicit features, 2. the multi-scale strategy, and 3. the specific architectural components of the ResGrouped-MLP network. The results are analyzed on the combined ALL dataset to ensure statistical reliability.
  • Impact of Multi-modal Implicit Features
We first investigate the impact of different basis feature functions by evaluating the performance of MS-ISSM using only single feature types: Luma, Chroma, and Curvature. As illustrated in Fig. 7, while single features, particularly Luma, can provide reasonable quality estimations, they exhibit limitations in capturing the full spectrum of perceptual distortions. For instance, geometric distortions are less perceptible in pure color metrics, and vice versa. The fusion of multi-modal features yields superior robustness, confirming that combining geometry and color attributes is essential for aligning with the HVS.
  • Impact of Multi-scale Strategy
To verify the necessity of the multi-scale hierarchy, we evaluated the performance at individual spatial scales (High, Medium, and Low). As reported in Table IV, although individual scales achieve decent correlations, their performance fluctuates across different datasets due to varying point cloud densities and content characteristics. For example, the High scale performs better on datasets with fine textures, while the Low scale is more effective for global structural distortions. By integrating all three scales, MS-ISSM effectively aggregates local details and global topology, achieving consistent and optimal performance across diverse datasets.
TABLE IV: The Performance comparison of different scales.
Scale Criteria ICIP WPC SJTU M-PCCD High PLCC 0.752 0.837 0.876 0.791 SROCC 0.705 0.842 0.856 0.829 KROCC 0.594 0.651 0.670 0.672 RMSE 0.216 0.376 0.175 0.151 Medium PLCC 0.812 0.827 0.869 0.911 SROCC 0.825 0.834 0.863 0.926 KROCC 0.648 0.637 0.673 0.767 RMSE 0.221 0.349 0.179 0.010 Low PLCC 0.879 0.814 0.788 0.855 SROCC 0.805 0.809 0.749 0.891 KROCC 0.631 0.612 0.555 0.722 RMSE 0.290 0.286 0.198 0.133
  • Impact of ResGrouped-MLP Architecture
Finally, to justify the design rationale of our regression network, we conducted an ablation study by removing or replacing key modules in the ResGrouped-MLP. The comparison results are summarized in Table V. w/o Log-Modulus Transformation: We removed the Log-Modulus preprocessing and used standard Z-score normalization directly on the raw coefficients. As shown in Table V, this resulted in a performance drop, with SROCC decreasing by 0.041. This confirms that the raw statistical features follow a heavy-tailed distribution, and the proposed Log-Modulus transformation effectively suppresses outliers, enabling the network to learn more robust feature representations. w/o Grouped Encoders (Early Concatenation): To validate the ”Split-Transform-Merge” strategy, we replaced the grouped encoders with a standard MLP that concatenates all multi-scale features at the input stage. The results show that the grouped strategy outperforms early concatenation by 0.03 in PLCC. This suggests that processing Luma, Chroma, and Curvature features independently in the early layers prevents information interference, allowing the network to capture distinct distortion patterns for each channel. w/o Attention Mechanism: We removed the Scale-wise Channel Attention blocks. The decline in performance indicates that the attention mechanism plays a crucial role. It allows the model to adaptively recalibrate the importance of different channels (e.g., assigning higher weights to Luma or Low-frequency geometry), thereby better mimicking the varying sensitivities of the HVS. In summary, each component of the MS-ISSM, from implicit feature extraction to the hierarchical regression network, makes a significant contribution to the final prediction accuracy and robustness.
TABLE V: ABLATION STUDY OF THE PROPOSED RESGROUPED-MLP ARCHITECTURE ON THE ”ALL” DATASET.
Model Variant PLCC SROCC KROCC RMSE Full MS-ISSM 0.913 0.914 0.751 0.188 w/o Log-Modulus 0.886 0.873 0.694 0.201 w/o Grouped Encoders 0.883 0.868 0.688 0.204 w/o Attention Block 0.901 0.904 0.717 0.192
  • The testing results of Cross-dataset.
We conducted cross-dataset evaluations to further validate the effectiveness and generalization capability of the proposed method. Given that the ICIP dataset is significantly smaller than the other datasets, it was used as the test set. We trained the model on the SJTU, M-PCCD, and WPC datasets, respectively, and tested the performance of MS-PSSIM on ICIP. As shown by the experimental results in Table VI, models trained on a single dataset generally performed well when validated on the ICIP dataset. These results demonstrate that the proposed method exhibits strong generalization ability and effectiveness.
TABLE VI: The cross-dataset evaluation results (Test Set: ICIP).
Training Set Criteria PLCC SROCC KROCC RMSE WPC 0.818 0.814 0.643 0.195 SJTU 0.880 0.878 0.694 0.159 M-PCCD 0.819 0.838 0.653 0.199
VI Conclusion
This paper presents a multi-scale implicit structural similarity (MS-ISSM) method for point cloud quality assessment (PCQA). To avoid the accumulation of matching errors in unstructured point clouds, the method leverages implicit functions to represent multi-scale features and evaluates quality based on differences in their coefficients. A ResGrouped-MLP network is introduced, incorporating a Log-Modulus transformation that stabilizes gradient descent and accelerates convergence. The architecture employs a grouped encoding strategy combined with Residual Blocks and Channel-wise Attention, enabling the model to preserve distinct physical semantics of luma, chroma, and geometry while adaptively highlighting the most salient distortions across high, medium, and low scales. Experiments demonstrate that MS-ISSM outperforms existing PCQA metrics on public datasets, providing a reliable and consistent quality evaluation.